The pressure of the oxygen gas inside a canister with a fixed volume is 5.0atm at 15oC. What is the pressure of the oxygen gas inside the canister if the temperature changes to 263K? Assume the amount of gas remains constant.
1
Expert's answer
2020-12-17T07:38:55-0500

T1 = 15 + 273.15 = 288.15 K

P1/T1 = P2/T2

P2 = P1T2/T1

P2 = 5.0 x 263 / 288.15 = 4.57 atm